# On Strongly Regular Graphs and the Friendship Theorem

### v1

@@ 3 The Friendship Theorem: An Alternative Proof
Let $\mathsf{G}$ be a finite graph in which any two distinct vertices have a single common neighbor. Then, $\mathsf{G}$ has a vertex that is adjacent to every other vertex.
@@
Remark 3.1 (On Theorem 3.1).
The windmill graph (see Figure 1) has the desired property of the friendship theorem and turns out to be the only graph with that property. Remarkably, the friendship theorem does not hold for infinite graphs. Indeed, for an inductive construction of a counterexample, one may start with a 5-cycle $\operatorname{\mathsf{C}}_{5}$, and repeatedly add a common neighbor for every pair of vertices that does not yet have one. This process results in a countably infinite friendship graph without a vertex adjacent to all other vertices.
Next, we present an alternative proof of Theorem 3.1, based on our expression in Corollary 2.1 for the Lovász $\vartheta$-function of strongly regular graphs and their complements, which are also strongly regular.
@@
Suppose the assertion is false, and let $\mathsf{G}$ be a counterexample — a finite graph in which any two distinct vertices have a single common neighbor, yet no vertex in $\mathsf{G}$ is adjacent to all other vertices. A contradiction is obtained as follows:
The first step shows that the graph $\mathsf{G}$ is regular, as proved in Chapter 44 of [1]. We provide a variation of that regularity proof, and then take a different approach, relying on graph invariants such as $\operatorname{\omega}(\mathsf{G}),\operatorname{\chi}(\mathsf{G})$, and $\vartheta(\mathsf{G})$.
@@
  • The given hypothesis yields that $\mathsf{G}$ is a connected graph. Let $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$, and let $\mathcal{N}(u)$ and $\mathcal{N}(v)$ denote, respectively, the sets of neighbors of the nonadjacent vertices $u$ and $v$.
@@
  • By symmetry, swapping $u$ and $v$ (as nonadjacent vertices) also yields $|\mathcal{N}(v)|\leq|\mathcal{N}(v)|$, so we get $d(u)=|\mathcal{N}(u)|=|\mathcal{N}(v)|=d(v)$ for all vertices $u,v\in\operatorname{\mathsf{V}}(\mathsf{G})$ such that $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$.
  • To complete the proof that $\mathsf{G}$ is regular, let $u$ and $v$ be nonadjacent vertices in $\mathsf{G}$. By assumption, except of one vertex, all vertices are either nonadjacent to $u$ or $v$. Hence, except of that vertex, all these vertices must have identical degrees by what we already proved.
  • Finally, by our further assumption (later leading to a contradiction), since there is no vertex in $\mathsf{G}$ that is adjacent to all other vertices, also the single vertex that is adjacent to $u$ and $v$ has a nonneighbor in $\mathsf{G}$, so it also should have an identical degree to all the degrees of the other vertices by what is proved above. Consequently, $\mathsf{G}$ is a regular graph.
@@
  • Let $\mathsf{G}$ be a $k$-regular graph on $n$ vertices. By assumption, every two vertices have exactly one common neighbor, so $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$.
  • First, if $k=1$ or $k=2$, then by assumption, it follows that $\mathsf{G}=\operatorname{\mathsf{K}}_{1}$ or $\mathsf{G}=\operatorname{\mathsf{K}}_{2}$, respectively, leading to a contradiction. Hence, let $k\geq 3$.
  • Two adjacent vertices in $\mathsf{G}$ share a common neighbor, so $\mathsf{G}$ contains a triangle. Moreover, $\mathsf{G}$ is $\operatorname{\mathsf{C}}_{4}$-free since every two vertices have exactly one common neighbor, so it is $\operatorname{\mathsf{K}}_{4}$-free. Hence, $\operatorname{\omega}(\mathsf{G})=3$.
  • We next show that $\operatorname{\chi}(\mathsf{G})=3$. First, $\operatorname{\chi}(\mathsf{G})\geq\operatorname{\omega}(\mathsf{G})=3$. We also need to show that $\operatorname{\chi}(\mathsf{G})\leq 3$, which means that three colors suffice to color the vertices of $\mathsf{G}$ in a way that no two adjacent vertices are assigned the same color. This can be done recursively by observing that each edge belongs to exactly one triangle (otherwise, the endpoints of that edge would have more than one common neighbor, which is not allowed). Moreover, each newly colored vertex always completes a properly colored triangle, ensuring that the coloring remains valid at every step without requiring a fourth color.
  • By the sandwich theorem $\operatorname{\omega}(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\operatorname{\chi}(\mathsf{G})$, so $\operatorname{\omega}(\mathsf{G})=\operatorname{\chi}(\mathsf{G})=3$ implies that $\vartheta(\overline{\mathsf{G}})=3$.
  • By Corollary 2.1, where $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$, we also get
    $\vartheta(\overline{\mathsf{G}})=1+\frac{k}{\sqrt{k-1}}.$
@@
    which completes the proof of the theorem.
A Second Alternative Proof of Theorem 3.1: From the point where we get, by contradiction, that $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$, it is possible to get a contradiction in the following alternative way. This second proof forms a variation of the proof presented in [1, Chapter 44].
@@
Let $\mathsf{G}$ be a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the following holds:
@@
    $\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}$ is an integer whose absolute value is at most $n-1$.
@@
This holds by relying on some basic properties of strongly regular graphs [2] as follows.
@@
  • Item 2 holds by the integrality and nonnegativity of the multiplicities $m_{1,2}$ of the second-largest and least eigenvalues of the adjacency matrix, which are respectively given by
@@
  • By applying Item 1 in Proposition 3.1 to $\mathsf{G}$ with the parameters $d=k$ and $\lambda=\mu=1$, we get $n=k^{2}-k+1$. This does not lead to a contradiction since summing over all the degrees of the neighbors of an arbitrary vertex $u$ in $\mathsf{G}$ gives $k^{2}$. Then, by the assumption of the theorem that every two vertices have exactly one common neighbor, it follows that the above summation counts every vertex in $\mathsf{G}$ exactly one time, except of $u$ that is counted $k$ times. Hence, indeed $n=k^{2}-k+1$.
  • By Item 2 in Proposition 3.1 with $d=k$ and $\lambda=\mu=1$, we get that $\frac{k}{\sqrt{k-1}}\in\mathbb{N}$. Consequently, $(k-1)|k^{2}\in\mathbb{N}$. Since $k^{2}=(k-1)(k+1)+1$, it follows that $(k-1)|1$, so $k=2$. If $k=2$, the only graph that satisfies the condition of Theorem 3.1 is $\mathsf{G}=\operatorname{\mathsf{K}}_{2}$, which also satisfies the assertion of the theorem. Hence, this argument contradicts the assumption in the proof since it led to the conclusion that $\mathsf{G}$ is a strongly regular graph within the family $\operatorname{\mathsf{srg}}(n,k,1,1)$.

### v3

@@ 3.1 The friendship theorem
Let $\mathsf{G}$ be a finite graph in which any two distinct vertices have a single common neighbor. Then, $\mathsf{G}$ consists of edge-disjoint triangles sharing a single vertex that is adjacent to every other vertex.
@@
Remark 3.1 (On Theorem 3.1).
The windmill graph (see Figure 1) has the desired property of the friendship theorem and it turns out to be the only graph with that property. Notably, the friendship theorem does not hold for infinite graphs. Indeed, for an inductive construction of a counterexample, one may start with a 5-cycle $\operatorname{\mathsf{C}}_{5}$, and repeatedly add a common neighbor for every pair of vertices that does not yet have one. This process results in a countably infinite friendship graph without a vertex adjacent to all other vertices.
@@
3.2 A new alternative proof of the friendship theorem relying on the Lovász $\vartheta$-function
This subsection presents an alternative proof of Theorem 3.1, utilizing the Lovász $\vartheta$-function of strongly regular graphs (and their complements) in Theorem 2.2. Recall that, by Proposition 2.1, the complements of strongly regular graphs are strongly regular as well.
@@ 3.2 A new alternative proof of the friendship theorem relying on the Lovász $\vartheta$-function
Suppose the assertion is false, and let $\mathsf{G}$ be a counterexample – a finite graph in which any two distinct vertices have a single common neighbor, yet no vertex in $\mathsf{G}$ is adjacent to all other vertices. A contradiction is obtained as follows:
The first step shows that the graph $\mathsf{G}$ is regular, as proved in Chapter 44 of [2]. We provide a variation of that regularity proof, and then take a different approach, relying on graph invariants such as $\omega(\mathsf{G}),\chi(\mathsf{G})$, and $\vartheta(\overline{\mathsf{G}})$.
@@
  • The given hypothesis yields that $\mathsf{G}$ is a connected graph (having a diameter of at most 2). Let $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$, and let $\mathcal{N}(u)$ and $\mathcal{N}(v)$ denote, respectively, the sets of neighbors of the nonadjacent vertices $u$ and $v$.
@@
  • By symmetry, swapping $u$ and $v$ (as nonadjacent vertices of the undirected graph $\mathsf{G}$) also yields $|\mathcal{N}(v)|\leq|\mathcal{N}(v)|$, so we get $d(u)=|\mathcal{N}(u)|=|\mathcal{N}(v)|=d(v)$ for all vertices $u,v\in\operatorname{\mathsf{V}}(\mathsf{G})$ such that $\{u,v\}\not\in\operatorname{\mathsf{E}}(\mathsf{G})$.
  • To complete the proof that $\mathsf{G}$ is regular, let $u$ and $v$ be fixed, nonadjacent vertices in $\mathsf{G}$. Consequently, $d(u)=d(v)$. By the assumption of the theorem, except for one vertex, all other vertices are either nonadjacent to $u$ or $v$. Hence, except for that vertex, all these vertices must have identical degrees by what we already proved.
  • Finally, by our further assumption (later leading to a contradiction), since there is no vertex in $\mathsf{G}$ that is adjacent to all other vertices, also the single vertex that is adjacent to $u$ and $v$ has a nonneighbor in $\mathsf{G}$, so it also should have an identical degree to all the degrees of the other vertices (by what is proved in the previous item). Consequently, $\mathsf{G}$ is a regular graph.
@@
  • Let $\mathsf{G}$ be a $k$-regular graph on $n$ vertices. By the theorem’s assumption, every two distinct vertices have exactly one common neighbor, so $\mathsf{G}$ is either a connected, strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,k,1,1)$ if $k\geq 3$, or $\mathsf{G}=\operatorname{\mathsf{K}}_{3}$ if $k=2$, or $\mathsf{G}=\operatorname{\mathsf{K}}_{1}$ if $k=0$ (recall that complete graphs are excluded from the family of strongly regular graphs).
  • First, $k=0$ and $k=2$ are excluded since complete graphs do not contradict our assumption of having a vertex not adjacent to all other vertices. Hence, let $k\geq 3$.
  • Every two adjacent vertices in $\mathsf{G}$ share a common neighbor, so $\mathsf{G}$ contains a triangle. Moreover, $\mathsf{G}$ is $\operatorname{\mathsf{C}}_{4}$-free since every two vertices have exactly one common neighbor, so it must be also $\operatorname{\mathsf{K}}_{4}$-free. Hence, $\omega(\mathsf{G})=3$.
  • We next show that $\chi(\mathsf{G})=3$. First, $\chi(\mathsf{G})\geq\omega(\mathsf{G})=3$. We also need to show that $\chi(\mathsf{G})\leq 3$, which means that three colors suffice to color all the vertices of $\mathsf{G}$ in a way that no two adjacent vertices are assigned the same color. This can be done recursively by observing that each edge belongs to exactly one triangle (otherwise, the endpoints of that edge would have more than one common neighbor, which is not allowed). Moreover, each newly colored vertex always completes a properly colored triangle, ensuring that the coloring remains valid at every step of the recursion without requiring a fourth color.
  • By the sandwich theorem $\omega(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\chi(\mathsf{G})$, so $\omega(\mathsf{G})=\chi(\mathsf{G})=3$ implies that $\vartheta(\overline{\mathsf{G}})=3$.
  • By Theorem 2.2, where $\mathsf{G}$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,k,1,1)$, we also get
    $\vartheta(\overline{\mathsf{G}})=1+\frac{k}{\sqrt{k-1}}.$ (19)
@@
    which completes the proof of the theorem by contradiction. Note that every edge in $\mathsf{G}$ lies on a triangle (by the theorem’s assumption), and $\mathsf{G}$ consists of edge-disjoint triangles since it is $\operatorname{\mathsf{C}}_{4}$-free.
3.3 Another proof of the friendship theorem relying on spectral properties of strongly regular graphs
This subsection presents an alternative proof of the friendship theorem, relying on the adjacency spectrum of strongly regular graphs. This second proof forms a variation of the proof provided in Chapter 44 of [2].
From the point in Section 3.2 where we get, by contradiction, that $\mathsf{G}$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,k,1,1)$ with $k\geq 3$, it is possible to get a contradiction in an alternative way, which relies on the following known result for strongly regular graphs.
@@ 3.3 Another proof of the friendship theorem relying on spectral properties of strongly regular graphs
Let $\mathsf{G}$ be a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the following holds:
@@
    $n-1-\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}$ is a nonnegative even integer.
@@
This holds by relying on some basic properties of strongly regular graphs (see Chapter 21 of [11]) as follows.
@@
  • Item 2 holds by the integrality and nonnegativity of the multiplicities $m_{1,2}$ of the second-largest and least eigenvalues of the adjacency matrix, respectively, which satisfy
    $0\leq 2m_{1,2}=n-1\mp\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}.$ (21)
@@
  • By applying Item 1 in Proposition 3.1 to $\mathsf{G}$ with the parameters $d=k$ and $\lambda=\mu=1$, we get $n=k^{2}-k+1$. This does not lead yet to a contradiction since summing over all the degrees of the neighbors of an arbitrary vertex $u$ in $\mathsf{G}$ gives $k^{2}$. Then, by the assumption of the theorem that every two vertices have exactly one common neighbor, it follows that the above summation counts each vertex in $\mathsf{G}$ exactly once, except for vertex $u$ that is counted $k$ times. Hence, indeed $n=k^{2}-k+1$.
[…]
3.4 Further discussions and insights on extensions of the friendship theorem
This subsection considers further generalizations and provides discussions on the friendship theorem and its proofs. The following discussions are presented in Remarks 3.23.4 and are supplemented by examples that examine strongly regular graphs of different orders, both small and large. In particular, Remark 3.4 builds on our proof technique, which is based on the Lovász $\vartheta$-function of strongly regular graphs, to offer insights into the extended friendship theorem studied in [3]. It also establishes additional results, which are further illustrated through Examples 3.13.3.
Remark 3.2.
By the frienship theorem (Theorem 3.1), every finite graph in which each pair of distinct vertices shares exactly one common neighbor must be isomorphic to a windmill graph (see Figure 1). Consequently, its order $n$ must be an odd integer, and the size of the graph is given by $|\operatorname{\mathsf{E}}(\mathsf{G})|=\tfrac{3}{2}(n-1)$. By definition, a graph $\mathsf{G}$ satisfying the assumption of Theorem 3.1 is $\operatorname{\mathsf{C}}_{4}$-free. A combinatorial proof, based on double counting, asserts that the size of a $\operatorname{\mathsf{C}}_{4}$-free graph $\mathsf{G}$ on $n$ vertices satisfies (see pp. 200–201 in [2])
$|\operatorname{\mathsf{E}}(\mathsf{G})|\leq\Bigl{\lfloor}\tfrac{1}{4}n(1+\sqrt{4n-3})\Bigr{\rfloor}.$ (22)
Furthermore, it has been shown that this upper bound on the size of $\mathsf{G}$ is nearly tight in the sense that if $p$ is a prime number, and $n=p^{2}+p+1$, then a specific construction of a $\operatorname{\mathsf{C}}_{4}$-free graph on $n$ vertices has a size that is given by (see pp. 201–202 in [2])
$|\operatorname{\mathsf{E}}(\mathsf{G})|=\tfrac{1}{4}(n-1)(1+\sqrt{4n-3}).$ (23)
The significant difference between the size of a windmill graph on $n$ vertices, scaling linearly as $\tfrac{3}{2}n$, and the largest size of a general $\operatorname{\mathsf{C}}_{4}$-free graph on $n$ vertices, scaling super-linearly as $\tfrac{1}{2}n^{\frac{3}{2}}$, stems from additional structural constraint on $\mathsf{G}$. Specifically, beyond being a $\operatorname{\mathsf{C}}_{4}$-free graph, $\mathsf{G}$ must satisfy a stronger condition that uniquely determines it (up to isomorphism) as a windmill graph.
Remark 3.3.
The friendship theorem was recently generalized in [12] to directed graphs (digraphs), allowing for asymmetry in liking relationships. The digraph version of the friendship theorem (see Theorem 1.1 of [12]) states that if a finite, simple, directed graph has the property that every pair of vertices has exactly one common out-neighbor, then it must be either:
  1. (1)
    A $k$-regular digraph – where each vertex has out-degree $k$ and in-degree $k$ – on $n=k^{2}-k+1$ vertices with $k\geq 2$, or
  2. (2)
    A so-called ”fancy wheel digraph,” which consists of a disjoint union of directed cycles with an additional vertex that has arcs to and from every vertex on these cycles.
Notably, the order $n$ of the $k$-diregular digraph matches that of the undirected graph $\mathsf{G}$ in the second proof (Section 3.3), in which it was established – prior to reaching the contradiction – that $\mathsf{G}$ is a $k$-regular graph of order $n=k^{2}-k+2$.
Remark 3.4.
As a possible extension of the friendship theorem, Proposition 1 of [3] states that if $\mathsf{G}$ is a finite, simple, and undirected graph such that every pair of vertices has exactly $\ell$ common neighbors for some fixed $\ell\geq 2$, then $\mathsf{G}$ is a regular graph. Consequently, $\mathsf{G}$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\ell,\ell)$. Furthermore, by Item 1 of Proposition 3.1 (here), it follows by substituting $\lambda=\mu=\ell$ that the parameters $n,d,\ell$ are related by the equation $(n-1)\ell=d(d-1)$. As an easy generalization of (19) in our proof of the friendship theorem, it follows from Theorem 2.2 that
$\vartheta(\overline{\mathsf{G}})=1+\frac{d}{\sqrt{d-\ell}}.$ (24)
Similar to our proof of the friendship theorem, which is based on the sandwich theorem stating that $\omega(\mathsf{G})\leq\vartheta(\overline{\mathsf{G}})\leq\chi(\mathsf{G})$, it follows that the clique and chromatic numbers of $\mathsf{G}$ satisfy, respectively,
$\omega(\mathsf{G})\leq 1+\biggl{\lfloor}\frac{d}{\sqrt{d-\ell}}\biggr{\rfloor},$ (25)
$\chi(\mathsf{G})\geq 1+\biggl{\lceil}\frac{d}{\sqrt{d-\ell}}\biggr{\rceil}.$ (26)
Furthermore, by Corollary 2.1, it follows from (24) that
$\vartheta(\mathsf{G})=\frac{n\sqrt{d-\ell}}{d+\sqrt{d-\ell}}.$ (27)
Similarly, by the sandwich theorem, the independence number of $\mathsf{G}$ and the chromatic number of its complement $\overline{\mathsf{G}}$ satisfy the following bounds:
$\alpha(\mathsf{G})\leq\biggl{\lfloor}\frac{n\sqrt{d-\ell}}{d+\sqrt{d-\ell}}\biggr{\rfloor},$ (28)
$\chi(\overline{\mathsf{G}})\geq\biggl{\lceil}\frac{n\sqrt{d-\ell}}{d+\sqrt{d-\ell}}\biggr{\rceil}.$ (29)
Hence, our proposed proof of the friendship theorem that relies on the Lovász $\vartheta$-function of strongly regular graphs, combined with Theorem 2.2 and Corollary 2.1 (here), and Proposition 1 of [3], lead to bounds on graph invariants of $\mathsf{G}$ and $\overline{\mathsf{G}}$ with respect to the extended version of the friendship theorem of [3] for an arbitrary $\ell\geq 2$.
The next three examples, preceded by definitions and notation, refer to Remark 3.4.
Definition 3.1 (Cartesian product of graphs).
Let $\mathsf{G}$ and $\mathsf{H}$ be graphs. The Cartesian product, denoted by $\mathsf{G}\,\square\,\mathsf{H}$, is a graph with a vertex set that is given by $\operatorname{\mathsf{V}}(\mathsf{G}\,\square\,\mathsf{H})=\operatorname{\mathsf{V}}(\mathsf{G})\times\operatorname{\mathsf{V}}(\mathsf{H})$ (i.e., it is equal to the Cartesian product of the vertex sets of $\mathsf{G}$ and $\mathsf{H}$), and its edge set is characterized by the property that distinct vertices $(g,h)$ and $(g^{\prime},h^{\prime})$, where $g,g^{\prime}\in\operatorname{\mathsf{V}}(\mathsf{G})$ and $h,h^{\prime}\in\operatorname{\mathsf{V}}(\mathsf{H})$, are adjacent in $\mathsf{G}\,\square\,\mathsf{H}$ if and only if one of the following conditions holds:
(1) $g=g^{\prime}$ and $\{h,h^{\prime}\}\in\operatorname{\mathsf{E}}(\mathsf{H})$, or   (2) $\{g,g^{\prime}\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ and $h=h^{\prime}$.
Definition 3.2 (Line Graph).
Let $\mathsf{G}=(\operatorname{\mathsf{V}},\operatorname{\mathsf{E}})$ be a graph. The line graph of $\mathsf{G}$, denoted by $\mathrm{L}(\mathsf{G})$, is a graph whose vertices are the edges of $\mathsf{G}$, and two vertices are adjacent in the line graph $\mathrm{L}(\mathsf{G})$ if the corresponding edges are incident in $\mathsf{G}$.
Example 3.1.
The Shrikhande graph and the Cartesian product of two complete graphs on 4 vertices, $\operatorname{\mathsf{K}}_{4}\square\operatorname{\mathsf{K}}_{4}$, are the two nonisomorphic strongly regular graphs in the family $\operatorname{\mathsf{srg}}(16,6,2,2)$, where each pair of distinct vertices has exactly two common neighbors. Note that the parameters $n=16$, $d=6$, and $\ell=2$ indeed satisfy the equality $(n-1)\ell=d(d-1)$. Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ denote, respectively, these two graphs. By (25), (26), (28), and (29), we get
$\alpha(\mathsf{G})\leq 4,\quad\omega(\mathsf{G})\leq 4,\quad\chi(\mathsf{G})\geq 4,\quad\chi(\overline{\mathsf{G}})\geq 4,$ (30)
which hold for $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$. A comparison with their exact values (computed by [13]) gives
$\alpha(\mathsf{G}_{1})=4,\quad\omega(\mathsf{G}_{1})=3,\quad\chi(\mathsf{G}_{1})=4,\quad\chi(\overline{\mathsf{G}}_{1})=6,$ (31)
$\alpha(\mathsf{G}_{2})=4,\quad\omega(\mathsf{G}_{2})=4,\quad\chi(\mathsf{G}_{2})=4,\quad\chi(\overline{\mathsf{G}}_{2})=4,$ (32)
so the four bounds on the graph invariants in (30) are tight for $\mathsf{G}=\mathsf{G}_{2}$.
Example 3.2.
For $n\geq 4$, the line graph of the complete graph on $n$ vertices, $\mathsf{G}=\mathrm{L}(\operatorname{\mathsf{K}}_{n})$, is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(\tfrac{1}{2}n(n-1),2(n-2),n-2,4)$. Consequently, the line graph $\mathrm{L}(\operatorname{\mathsf{K}}_{6})$ is in the family $\operatorname{\mathsf{srg}}(15,8,4,4)$, so every pair of distinct vertices in that graph has exactly four common neighbors. Using the SageMath software [13] gives $\alpha(\mathsf{G})=3$, $\omega(\mathsf{G})=5$, $\chi(\mathsf{G})=5$, and $\chi(\overline{\mathsf{G}})=4$. Substituting the graph parameters $n=15$, $d=8$, and $\ell=4$ into (25)–(29) shows that the three bounds in (25), (26), and (28) are tight. However, the bound in (29) is not tight as it gives $\chi(\overline{\mathsf{G}})\geq 3$, whereas its exact value is 4.
Example 3.3 (Complements of symplectic polar graphs).
The symplectic polar graphs are strongly regular graphs, belonging to the families $\operatorname{\mathsf{srg}}(v,k,\lambda,\mu)$, where
$v=\frac{q^{2n}-1}{q-1},\quad k=\frac{q(q^{2n-2}-1)}{q-1},\quad\lambda=\frac{q^{2n-2}-1}{q-1}-2,\quad\mu=\frac{q^{2n-2}-1}{q-1},$ (33)
for all $n\in\mathbb{N}$ and $q\geq 2$ that is a prime power. Consequently, $\lambda=\mu-2$ and $\mu=\frac{k}{q}$. That symplectic polar graph is denoted by $\mathsf{G}=\operatorname{\mathsf{Sp}}(2n,q)$ (see Section 2.5 in [9]). The complement graph of a strongly regular graph in the family $\operatorname{\mathsf{srg}}(\nu,k,\lambda,\mu)$ is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(v,v-k-1,v-2k+\mu-2,v-2k+\lambda)$ (see Proposition 2.1). Substituting $v,k,\lambda,\mu$ in (33) gives that $\mathsf{H}=\overline{\mathsf{G}}$ is a strongly regular graph in the family
$\vspace*{0.1cm}\operatorname{\mathsf{srg}}\,\Biggl{(}\frac{q^{2n}-1}{q-1},q^{2n-1},q^{2n-2}(q-1),q^{2n-2}(q-1)\Biggr{)}.$ (34)
Consequently, the family of the complements of the symplectic polar graphs, $\mathsf{H}=\overline{\operatorname{\mathsf{Sp}}(2n,q)}$, where $n\in\mathbb{N}$ and $q\geq 2$ is a prime power, forms an infinite family of strongly regular graphs which are characterized by the property that every pair of vertices in such a graph has an identical number of common neighbors, whose value is given by $\ell=q^{2n-2}(q-1)$.
Table I: (Example 3.3) Complements of symplectic polar graphs, the families of strongly regular graphs to which they belong (see (34)), the fixed number ($\ell$) of common neighbors for each pair of distinct vertices, the values of the Lovász $\vartheta$-function for these graphs and their complements, and their exact matchings with the independence and chromatic numbers of these graphs.
[…]
$3$ $2$ $\operatorname{\mathsf{srg}}(63,32,16,16)$ $16$ $7$ $7$ $9$ $9$
$3$ $3$ $\operatorname{\mathsf{srg}}(364,243,162,162)$ $162$ $13$ $13$ $28$ $28$
$3$ $4$ $\operatorname{\mathsf{srg}}(1365,1024,768,768)$ $768$ $21$ $21$ $65$ $65$
$3$ $5$ $\operatorname{\mathsf{srg}}(3906,3125,2500,2500)$ $2500$ $31$ $31$ $126$ $126$
$3$ $7$ $\operatorname{\mathsf{srg}}(19608,16807,14406,14406)$ $14406$ $57$ $57$ $344$ $344$
$4$ $2$ $\operatorname{\mathsf{srg}}(255,128,64,64)$ $64$ $15$ $15$ $17$ $17$
$4$ $3$ $\operatorname{\mathsf{srg}}(3280,2187,1458,1458)$ $1458$ $40$ $40$ $82$ $82$
$4$ $4$ $\operatorname{\mathsf{srg}}(21845,16384,12288,12288)$ $12288$ $85$ $85$ $257$ $257$
$5$ $2$ $\operatorname{\mathsf{srg}}(1023,512,256,256)$ $256$ $31$ $31$ $33$ $33$
$5$ $3$ $\operatorname{\mathsf{srg}}(29524,19683,13122,13122)$ $13122$ $121$ $121$ $244$ $244$
$5$ $4$ $\operatorname{\mathsf{srg}}(349525,262144,196608,196608)$ $196608$ $341$ $341$ $1025$ $1025$
$6$ $2$ $\operatorname{\mathsf{srg}}(4095,2048,1024,1024)$ $1024$ $63$ $63$ $65$ $65$
$6$ $3$ $\operatorname{\mathsf{srg}}(265720,177147,118098,118098)$ $118098$ $364$ $364$ $730$ $730$
$7$ $2$ $\operatorname{\mathsf{srg}}(16383,8192,4096,4096)$ $4096$ $127$ $127$ $129$ $129$
$8$ $2$ $\operatorname{\mathsf{srg}}(65535,32768,16384,16384)$ $16384$ $255$ $255$ $257$ $257$
$9$ $2$ $\operatorname{\mathsf{srg}}(262143,131072,65536,65536)$ $65536$ $511$ $511$ $513$ $513$
In regard to their graph invariants (as illustrated in Table I), by Section 2.5.4 of [9],
$\alpha(\mathsf{H})=\omega(\mathsf{G})=\frac{q^{n}-1}{q-1},$ (35)
and, by Theorem 3.29 of [14],
$\vartheta(\mathsf{H})=\alpha(\mathsf{H})=\frac{q^{n}-1}{q-1},$ (36)
$\vartheta(\overline{\mathsf{H}})=\chi(\mathsf{H})=q^{n}+1.$ (37)
Consequently, it follows that the bounds in (28) and (29), applied to $\mathsf{H}$ for its independence and chromatic numbers, are tight. Numerical results are presented in Table I.